# ATOMIUM: Dust and tracers of binarity in the continua

### v1

@@ 3.4.2 W Aql
W Aql has a known main sequence (F9) companion at a projected separation of , corresponding to a physical separation au (Danilovich et al., 2015a, 2024). The individual higher resolution 12m array continuum images, plotted in Fig. LABEL:fig:waql, show a secondary peak, detected with a certainty of , to the south west of the primary continuum peak. Taking the primary peak as the position of the AGB star, the secondary peak corresponds well to the location of the F9 companion, which was observed with VLT/SPHERE contemporaneously with the extended array data (9 July 2019, Montargès et al., 2023, a day after the ALMA observations were completed). We indicate the position of the companion as observed with SPHERE by the yellow cross in Fig. LABEL:fig:waql. While the primary peak is well-resolved with the extended array, the secondary peak appears as a point source in the extended and combined continuum maps. It is only marginally resolved from the primary peak in the mid continuum and not at all in the compact continuum (Fig. 11). At this wavelength (1.24 mm), the F9 companion is well below the detection threshold, so this significant detection indicates some interaction between the circumstellar AGB dust and the F9 star, which is discussed in more detail in Sect. 5.1.2.
@@ 3.4.4 Gru
Homan et al. (2020) previously presented the extended continuum map of Gru, which shows a close secondary peak to the south-west of the primary continuum peak. Here we present the continuum maps obtained from the other ALMA arrays and the combined continuum map for the first time (Figs. LABEL:fig:pi1gru, 11). We also present the ACA Band 3 continuum map (Fig. 12).
The secondary peak is marked by a black cross on the extended and combined continuum plots in Fig. LABEL:fig:pi1gru. We henceforth refer to this feature as the C companion, following Montargès et al. (2025). The AGB star, which we associate with the position of the primary continuum peak, is the A component and the B component is a long-known main sequence G0V star (Feast, 1953), which, in the Gaia DR3 epoch (2016.0), is separated from the AGB star by 2.68″ (Gaia Collaboration et al., 2021), corresponding to a projected separation of 440 au. The B component is outside of the plotted field of views of the extended and combined data, but we indicate its approximate relative position (calculated from the Gaia catalogue by assuming that the separation between A and B did not significantly change between the Gaia epoch and our ALMA observations) as a yellow cross on the plot of the mid continuum image (Fig. LABEL:fig:pi1gru) and the compact continuum image in Fig. 11. We note that the compact continuum emission is relatively featureless. In both the mid and compact images there are small regions of emission in the direction of the B companion but, although they are within contours, they are smaller than the synthetic beam, so may not be significant. However, their presence in two separate observations suggests that there may be some gravitational influence from the B companion (similar to what was seen at the larger-scales examined for W Aql, as discussed in Danilovich et al., 2024). We do not detect any emission above the noise near the position of the B companion, in contrast with W Aql, which may be because the companion is more distant and situated in a less dense wind (e.g. compare the mass-loss rates in Table 13).
@@ 3.4.5 AH Sco
Finally, we note that all the continuum maps of AH Sco have several small regions of flux scattered around the main region of continuum emission. These may represent dispersed clumps of dust in the circumstellar environment.
@@ 4 Analysis
4.1 Resolved-out flux
Because ALMA is an interferometer, it is not guaranteed to recover all the flux associated with a source, in particular it may resolve out some smooth large scale flux. The maximum recoverable scale (MRS) is the scale at which smooth regions of flux can be recovered with confidence. In general, the MRS is larger for the more compact (lower angular resolution) configurations of ALMA. The MRS for the main ATOMIUM dataset are – for the extended array data, for the mid array data, and 8– for the compact array data. More precise MRS for each star and configuration are given in Table E.2 of Gottlieb et al. (2022). The combined MRS is generally as good as for the compact configuration data, except for AH Sco and KW Sgr, which were not observed with the compact configuration, and for which the MRS is as good as for the mid configuration data. For the new ACA data presented here, the MRS range from 64–77″ and are given for each star in Table 9.
Four ATOMIUM stars are also in the DEATHSTAR survey (Ramstedt et al., 2020), which observed a large sample of AGB stars using the ALMA ACA in Bands 6 and 7. For their Band 6 observations, covering similar frequencies to ATOMIUM albeit over a smaller range, the MRS was (Ramstedt et al., 2020). We compared the self-calibrated Band 6 continuum fluxes from DEATHSTAR with those from our ATOMIUM combined and compact continuum images, for the overlapping stars (IRC10529, SV Aqr, T Mic and IRC+10011). We found no evidence of more flux being resolved out for the ATOMIUM data than for the DEATHSTAR data. However, this does not mean that all the flux has necessarily been recovered. Herschel/PACS images of dust at 70 and 160 m are available for some of the stars in our sample. For W Aql, Gru, R Hya, and T Mic, extended flux is seen on scales of from the position of the star, indicating significant dust at larger scales than can be recovered with ALMA (Cox et al., 2012). This is significant because at longer wavelengths (i.e. 1.2–1.3 mm for ALMA Band 6) we might detect cooler dust, which we would expect to be located further from the star. However, since ALMA cannot recover flux on such large scales — scales that can be larger than the ALMA field of view for a single pointing — we would not expect the majority of flux from such a large dust shell to be recovered, even when using the ACA. Comparing our total combined fluxes (Table 8) with single-dish observations from Altenhoff et al. (1994) and Dehaes et al. (2007), we can estimate the recovered fluxes for R Aql, R Hya, and IRC+10011 as around 60%, 70%, and 6–12%, respectively. The large amount of resolved out flux for IRC+10011 (88–94% lost) fits with the negative artefacts we see in Fig. LABEL:fig:irc+10011.
To give an indication of how much extended vs compact continuum emission is detected around each star, we compared the UD flux density (see Table 2 and Sect. 2.4) with the total flux recovered in our combined images (Table 8). We found no clear correlation with distance overall. Excluding the RSGs, we found a weak positive correlation between the flux ratio and distance, but not to a significant level. Most importantly, we see a large spread in the flux ratios, from 1.07 (U Del) to 2.24 (IRC10529). U Del having the lowest flux ratio fits with what we found from the UD-subtracted map (Fig. LABEL:fig:udel) which shows that the majority of the continuum flux is associated with a uniform disc of emission at the stellar position. The stars with the highest ratio (IRC10529 followed by W Aql) exhibit flux away from the continuum peak. Some of the other stars with lower flux ratios nevertheless have extended features, likely arising from dust, in their continuum maps, or are known to be dusty based on other observations (e.g. Herschel/PACS imaging Cox et al., 2012; Maercker et al., 2022). This discrepancy could be caused by resolved-out flux, amplitude errors (especially for sources with high dynamic ranges ) or a combination of factors. We also checked whether there was any correlation between the ratio of total flux and UD flux density with the apparent or physical UD radii and found no trends.
@@ 4.2 Spectral indices
For the seven stars in our ACA Band 3 sample we also calculate spectral indices. Because all the stars are unresolved in the large (10–15″) synthesised beams of the ACA data, we use the peak fluxes to approximate the total flux. We also fit 2D Gaussians to the observations to extract the integrated flux and found equivalent values, within the uncertainties. The peak fluxes are listed in Table 9. We adopt 104.96 GHz as the central frequency of the Band 3 ACA continua and 241.75 GHz as the central frequency of the Main Array Band 6 combined continua. The spectral indices are then calculated from these values, the ACA peak fluxes and the combined UD flux densities (Table 2). The resultant spectral indices are listed in Table 9 and fall in the range 1.5–2.5. Within our uncertainties, these agree with the unresolved spectral indices calculated by Bojnordi Arbab et al. (2024) from theoretical models of AGB stars.
We note that the 12-m array Band 6 measurements from the UD fits do not include scattered, extended emission. Since the ACA resolution is greater than the maximum Band 6 MRS, some of the emission detected by the ACA may be either resolved out, or undetected due to low surface brightness at Band 6. On the other hand, the rms noise in the ACA observations is around 0.2 mJy, whilst the 12-m combined rms is mostly mJy, so the ACA flux may be depressed by spatial decorrelation, especially as the S/N was too low for self-calibration. To summarise these effects, if the Band 6 data has more missing flux, then the spectral index would be underestimated, an effect that would be systematically worse for the nearby sources. If the Band 3 were decorrelated, the result would be a steeper spectral index, especially for the weaker sources, which are more likely to be more distant.
The stars in our sample with the largest spectral indices (, see Table 9) are R Aql, GY Aql and T Mic. We find that R Aql and T Mic have strong ACA continuum detections, in excess of , so are unlikely to suffer from decorrelation. GY Aql and IRC10529 have comparatively weak ACA continuum detections (5–) so decorrelation could explain its high spectral index.
@@
We checked the data for possible correlations between spectral index and other parameters. We find no significant correlation between spectral index and distance, suggesting that there is no factor uniformly affecting only the closest or most distant stars. We checked for correlations between the spectral index and the individual ACA or 12-m fluxes and found none. Finally, we also found no apparent correlation between spectral index and stellar pulsation period.
@@ 4.3 Comparison with NIR stellar radii
Star | Ref. for | Calc | Calc diameter | NIR | NIR diameter | Ref. for | ||
@@
where and mas. This relationship is plotted as the dashed line in the right panel of Fig. 3, with the uncertainties shown by they grey shaded region. Note that although IRC10529 and IRC+10011 are outliers with larger NIR diameters than ALMA UD diameters, the uncertainties on their NIR radii are so large that they do not significantly contribute to the fit. With the exception of these two stars, there is minimal scatter around the trend line for the stars with measured data. There is more scatter among the stars with calculated optical diameters, but the only true outlier among these is RW Sco, for which uncertainties come from being unresolved by ALMA (resulting in an uncertain UD diameter, see Table 2) and an uncertain effective temperature estimate, as noted above. Plotting only calculated sizes for all the stars would result in even more scatter in our plot. This can partly be explained by the variations in temperature and radius at different phases of the pulsation period. Bojnordi Arbab et al. (2024) also showed that the theoretical sizes of AGB stars can vary across pulsation period, which would add to the scatter given that observations of different stars and different configurations were taken at different pulsation phases (e.g. see Table 7 of Baudry et al., 2023). Given that the trend line from Eq. 2 does not pass through (0,0), we must assume that this relationship only holds over a certain (apparent) size regime, or is not truly linear. To better understand this, we require a larger sample of high quality measurements in both the NIR and at 1.24 mm, ideally observed contemporaneously.
@@ 4.4 Trends between stellar properties
We checked for any trends between the measured continuum properties, based on the UD fits. In Fig. 4 we plot the monochromatic luminosity, which is essentially a normalised UD flux density, against the measured UD radii for our sample. The monochromatic luminosity was calculated using
@@
No clear trend is seen for the SR AGB stars, although we note that all of the SR stars have radii cm ( au). They also all have monochromatic luminosities mJy pc2, whereas the Miras all have monochromatic luminosities mJy pc2. The dotted vertical and horizontal lines in Fig. 4 illustrate these boundaries between the Miras and SRs. We see a general trend for the Miras of increasing flux with increased radius. The RSGs are both larger and more luminous than the AGB stars, which is to be expected as they are also more massive.

### v2

@@ 3.5.3 W Aql
W Aql has a known main sequence (F9) companion at a projected separation of , corresponding to a physical separation au (Danilovich et al., 2015a, 2024). The individual higher resolution 12m array continuum images, plotted in Fig. LABEL:fig:waql, show a secondary peak, detected with a certainty of , to the south west of the primary continuum peak. Taking the primary peak as the position of the AGB star, the secondary peak corresponds well to the location of the F9 companion, which was observed with VLT/SPHERE contemporaneously with the extended array data (9 July 2019, Montargès et al., 2023, a day after the ALMA observations were completed). We indicate the position of the companion as observed with SPHERE by the yellow cross in Fig. LABEL:fig:waql. While some extended emission is resolved around the primary peak with the extended array, the secondary peak appears as a point source in the extended and combined continuum maps. It is only marginally resolved from the primary peak in the mid continuum and not at all in the compact continuum (Fig. 14). At this wavelength (1.24 mm), the F9 companion is well below the detection threshold, so this significant detection indicates a concentration of dust close to the F9 star, the origin of which is discussed in more detail in Sect. 5.1.2.
@@ 3.5.5 Gru
Homan et al. (2020) previously presented the extended continuum map of Gru, which shows a close secondary peak to the south-west of the primary continuum peak. Here we present the continuum maps obtained from the other ALMA arrays and the combined continuum map for the first time (Figs. LABEL:fig:pi1gru, 14). We also present the ACA Band 3 continuum map (Fig. 15).
The secondary peak is marked by a black cross on the extended and combined continuum plots in Fig. LABEL:fig:pi1gru. We henceforth refer to this feature as the C companion, following Montargès et al. (2025). The AGB star, which we associate with the position of the primary continuum peak, is the A component and the B component is a long-known main sequence G0V star (Feast, 1953), which, in the Gaia DR3 epoch (2016.0), is separated from the AGB star by 2.68″ (Gaia Collaboration et al., 2021), corresponding to a projected separation of 440 au. The B component is outside of the plotted field of views of the extended and combined data, but we indicate its approximate relative position (calculated from the Gaia catalogue by assuming that the separation between A and B did not significantly change between the Gaia epoch and our ALMA observations) as a yellow cross on the plot of the mid continuum image (Fig. LABEL:fig:pi1gru) and the compact continuum image in Fig. 14. We do not detect any emission above the noise near the position of the B companion, in contrast with W Aql, which may be because the companion is more distant and situated in a less dense wind (e.g. compare the mass-loss rates in Table 14).
@@ 4.1 Spectral indices
4.1.1 Intra-Band 6 spectral indices
Since our Band 6 observations covered frequencies spanning GHz, we were able to calculate intra-Band 6 spectral indices for our sample. In cleaning an image with tclean, as well as identifying the brightest pixels in each cycle, we used the option to fit for the linear dependence of flux with frequency, using the first two terms of a Taylor series, giving total intensity and spectral index (alpha) images. We use the alpha.error image produced by tclean to exclude unreliable spectral index measurements, but for a well-cleaned image, this is a lower limit to the actual error, since it is based only on the residual noise in the zeroth and first Taylor coefficient images. This does not include systematic errors such as in alignment or calibration (indeed, a well-known diagnostic for tiny spatial misalignments is a systematic gradient in alpha across an image).
Without additional calibrators/dedicated flux scale measurements, the accuracy of an individual ALMA observation at Band 6 is 5% at best. Each mid and extended configuration data set contains at least 4 observations at 4 sets of frequencies (see Gottlieb et al. 2022 Fig. 1 for the ranges covered in each tuning). Our calibration methods force the spectral index to be linear, but if both sets of the lowest and the highest frequencies have 5% errors in the opposite sense this could tilt the spectral index by about 0.3, regardless of S/N. Statistically, this is likely to affect one or two out of our 17 data sets, per configuration. Another factor is the potential presence of residual lines. The combined or extended configuration data possess the most baselines sampling the smallest angular scales, meaning that these can be used for measuring the spectral index of the stellar continuum. The S/N of any other continuum features (e.g. any extended emission) is much too low for a reliable spectral index measurement.
We find that for most stars, the uncertainty on the spectral index becomes very high further than beam from the continuum peak. Hence, to compute average spectral indices, , we only consider pixels (i) which are within the circular region corresponding to the major axis of the beam convolved with the size of the UD fit to the star and (ii) for which the uncertainty on the spectral index is . Then, for these selected pixels, we construct a histogram of the spectral indices and derive a Gaussian kernel density estimation (KDE). The peak of this KDE is taken to be the spectral index. The uncertainty is then conservatively taken to be the sum in quadrature of the full width at half maximum of the KDE and the maximum uncertainty on the individual pixels. To check the robustness of this estimate, we also considered the spectral index of the pixel closest to the continuum peak and the mean spectral index of all the selected pixels. Both methods agreed well with our final value.
Our calculated are give in Table 3. Example plots of the spectral index for W Aql and U Her are given in Fig. 4. The remaining stars are shown in Fig. 18 and the KDEs of the spectral indices are shown for all stars in Fig. 19. We were unable to extract a value of for RW Sco from the combined data because the low S/N resulted in very large uncertainties on the spectral index. However, using only the extended array data we were able to estimate a spectral index for RW Sco.
For Gru we get an anomalously high . This is higher than the Band 6 to Band 7 value, , found by Esseldeurs et al. (2025), and the Band 3 to Band 6 value we find, , details in Sect. 4.1.2. The Band 6 mid data gave a value closer to the multi-band measurements. We suspect that the extended data for this star was affected by all the possible sources of uncertainty mentioned above and thus acts as a caution in using in-band spectral indices.
| Star | Dust mass [] | ||
| GY Aql | |||
| R Aql | |||
| IRC10529 | |||
| W Aql | |||
| SV Aqr | |||
| U Del | |||
| Gru | |||
| U Her | |||
| R Hya | |||
| T Mic | |||
| S Pav | |||
| IRC+10011 | |||
| V PsA | |||
| RW Sco | |||
| KW Sgr | |||
| VX Sgr | |||
| AH Sco |
4.1.2 Band 3 to Band 6 spectral indices
@@
For the seven stars in our ACA Band 3 sample we also calculate band-to-band spectral indices, . Because all the stars are unresolved in the large (10–15″) synthesised beams of the ACA data, we use the peak fluxes to approximate the total flux. We also fit 2D Gaussians to the observations to extract the integrated flux and found equivalent values, within the uncertainties. The peak fluxes are listed in Table 11. We adopt 104.96 GHz as the central frequency of the Band 3 ACA continua and 238.45 GHz as the central frequency of the Main Array Band 6 compact configuration continua777Note that this is a slightly different value than for the combined array data because fewer spectral windows were observed with the compact configuration than with the extended and mid configurations; see Gottlieb et al. (2022) for details.. The spectral indices are then calculated from these values and the peak fluxes of the ACA and compact data (Table 9). The resultant spectral indices are listed in Table 3 and fall in the range 1.7–2.5, aside from RW Sco, for which . This is most likely the result of the low S/N measurements for this star.
4.1.3 Trends with spectral index
Within our uncertainties (aside from RW Sco and Gru), the and spectral indices agree with each other. Aside from the outliers already discussed, we find that the spectral indices for all our stars . This agrees with the theoretical predictions for unresolved spectral indices calculated by Bojnordi Arbab et al. (2024). Since we were only able to reliably derive the spectral indices close to the star (essentially only at the continuum peak), we would expect the main contribution to the spectral index to come from the star itself. This is borne out by the predictions of Bojnordi Arbab et al. (2024); although they find variations in the unresolved spectral index with pulsation phase, their values are in the range 1.7–2.05 over frequencies from 100–300 GHz, in good agreement with our results. It is likely that the values are also dominated by the stars since, for the same enclosed regions, the flux densities of the standard combined maps (Table 10) are a factor of 2–10 higher than the flux densities of the UD-subtracted maps (Table 12). In general, the value of of the spectral index is consistent with optically thick blackbody emission, another indication that the stellar contribution dominates.
@@ 4.1.3 Trends with spectral index
We checked the data for possible correlations between the spectral index and other parameters, searching for trends or systematics. We find no significant correlation between spectral index and distance, suggesting that there is no factor uniformly affecting only the closest or most distant stars. For we checked for correlations between the spectral index and the individual ACA or 12-m fluxes and found none. We also found no apparent correlation between spectral index and stellar pulsation period. There is a tentative positive correlation between and mass-loss rate (see Appendix G for details on the mass-loss rates), with low statistical significance. The models of Bojnordi Arbab et al. (2024) predict a slightly larger spread in unresolved spectral indices for their higher mass-loss rate model ( compared with their canonical model), but only to 1.55–2.1 in the 100–300 GHz range. We hence suggest that this tentative correlation between spectral index and mass-loss rate requires further observations to confirm.
4.2 Dust mass estimates
Having calculated the spectral indices for our sample, we can now estimate the mass of the dust visible in our ALMA observations. Following Knapp et al. (1993) we estimate the dust masses using
where is the total flux density of the UD-subtracted continuum image, is the radius of the dust grains, assumed to be spherical, which we take as 0.3 m, is the mass density of the dust grains, which we take as 3.3 g cm-3, is the speed of light, is the distance to the star (see Table 1), is the frequency, here 241.75 GHz, is the grain emissivity at , is the Boltzmann constant, and is the dust temperature, which we take to be the lower value out of the brightness temperature (Table 2) or 1000 K, which is close to the condensation temperature of silicate grains (Suh, 1999). We assume , where is the dust emissivity spectral index. We derive for each star from our values in Table 3, resulting in for most stars. For , we follow O’Gorman et al.’s (2015) modification of Knapp et al.’s (1993) estimate and use
Our resultant dust mass estimates are given in Table 3 and range from for SV Aqr to for AH Sco. These are all lower limits on the true dust masses around these stars because, as discussed in Sect. 3.1, a significant amount of flux has been resolved out for our observations. This is particularly apparent when comparing our results to dust masses calculated from Herschel/PACS observations by Cox et al. (2012). For the three stars in both samples, W Aql, R Hya, and T Mic, Cox et al. (2012) find dust masses of 1.6–, and 1–, respectively. This is three to four orders of magnitude larger than our dust masses, in all cases. Cox et al. (2012) assume rather low dust temperatures of 35 K, but even if we use this value in our calculations, we still find about two orders of magnitude less dust. We also investigated the impact of different values of on our dust masses, including the value of adopted by Cox et al. (2012) from Li & Draine (2001), and found that varying from to 2 resulted in less than a 35% change in the calculated dust mass. Finally, a more precise estimate of the dust mass would need to take into account non-spherical grains with a range of different sizes, which is beyond the scope of the present work, especially since emission from the larger grains, assumed to be found further from the star, is most likely to be resolved out.
@@ 4.3 Comparison with NIR stellar radii
| Star | Ref. for | Calc | Calc diameter | NIR diameter | Ref. for | |||
@@
where and mas. This relationship is plotted as the dashed line in the right panel of Fig. 5, with the uncertainties shown by the grey shaded region. Note that although IRC10529 and IRC+10011 are outliers with larger NIR diameters than ALMA UD diameters, the uncertainties on their NIR radii are so large that they do not significantly contribute to the fit. With the exception of these two stars, there is minimal scatter around the trend line for the stars with measured data. There is more scatter among the stars with calculated optical diameters, but the only true outlier among these is RW Sco, for which uncertainties come from being unresolved by ALMA (resulting in an uncertain UD diameter, see Table 2) and an uncertain effective temperature estimate, as noted above.
Of course Eq. 4 must only be valid over a limited regime, since a non-zero diameter at one wavelength cannot physically correspond to a diameter of zero in another. However, in the right panel of Fig. 5 we plot observed sizes not physical sizes and the real lower limits on both axes are the resolution limits. Therefore, we must assume that this relationship only holds over a certain (apparent) size regime, or is not truly linear. To better understand this, we require a larger sample of high quality measurements in both the NIR and at 1.24 mm, ideally observed contemporaneously.
If, instead of allowing a general linear fit, we force the line to pass through the origin, we find a slope of 1.38, which is still a reasonable fit to the measured data and is shown as the brown dot-dashed line in Fig. 5. Plotting only calculated sizes for all the stars would result in even more scatter in our plot. This can partly be explained by the variations in temperature and radius at different phases of the pulsation period. Bojnordi Arbab et al. (2024) also showed that the theoretical sizes of AGB stars can vary across pulsation period, which would add to the scatter given that observations of different stars and different configurations were taken at different pulsation phases (e.g. see Table 7 of Baudry et al., 2023).
Finally, in addition to examining the observed sizes of our stars, we tested fitting the physical sizes. This introduces additional uncertainties from the distances but also has the effect of rearranging the data so that less weight is placed on the closest stars which have the largest apparent sizes but similar physical sizes to the majority of our sample. Within the uncertainties, a fit to these parameters results in a 1:1 relation between the ALMA- and NIR-derived physical sizes. (This holds whether or not we include the RSGs and/or the two largest AGB stars, IRC10529 and IRC+10011, in our fit.) The results of this test indicate that the formal uncertainties on our fit (Eq. 3) are smaller than the real uncertainties.
@@ 4.4 Trends between stellar properties
We checked for any trends between the measured continuum properties, based on the UD fits. In Fig. 6 we plot the monochromatic luminosity, which is essentially a normalised UD flux density, against the measured UD radii for our sample. The 1.24 mm radio luminosity, which we refer to as the monochromatic luminosity, was calculated using
@@
No robust trend is seen for the SR AGB stars, although we note that all of the SR stars have radii cm ( au). They also all have monochromatic luminosities mJy pc2, whereas the Miras all have monochromatic luminosities mJy pc2. The dotted vertical and horizontal lines in Fig. 6 illustrate these boundaries between the Miras and SRs. We see a general trend for the Miras of increasing flux with increased radius. The RSGs are both larger and more luminous than the AGB stars, which is to be expected as they are also more massive.
Considering all the stars plotted in Fig. 6, there is a clear positive correlation between the UD radius and the monochromatic luminosity. This is to be expected since the UD radius is related to the physical radius of the star (as discussed in Sect. 4.3) and the monochromatic luminosity has a dependence on the bolometric luminosity. Since the bolometric luminosity is proportional to the square of the physical radius, see Eq. 3, it is then not surprising that a similar correlation would be reflected in the corresponding mm properties. A fit to these properties gives a slope close to 2, as expected for , but with large uncertainties owing to the scatter and uncertainties on our data.